OpenStudy (anonymous):

in a different 6th grade class the ratio of boys to girls is 3:2. how many boys and how many girls could there be in this class? is there more than one possible answer?

Vocaloid (vocaloid):

3:2 ratio means that there are 3 boys for every 2 girls so, the number of boys must be a multiple of 3 and the number of girls must be (2/5) times the number of boys

Vocaloid (vocaloid):

one example is 30 boys and 20 girls

Vocaloid (vocaloid):

*small correction, the number of girls must be (2/3) the number of boys

OpenStudy (texaschic101):

the ratio 3:2 can be written as 3/2. And any equivalent fraction would work. 3/2 = 6/4 = 9/6......so there can be 6 boys to 4 girls...or 9 boys to 6 girls...and so on.